OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian?

OpenStudy (anonymous):

hhelp

OpenStudy (anonymous):

Let v= the vegetarian wraps. Let c= the chicken wraps. Since the vegetarian wraps sell for $1.00, the money earned by selling vegetarian wraps is $1.00v. Since the chicken wraps sell for $1.80. the money earned by selling chicken wraps is $1.80c. So now you have two equations. Equation one refers to the numbed of wraps sold by type and total amount of wraps sold. So it looks like this: v+c=70 The other equation refers to the money each wrap earns and the total amount earned from selling both wraps. So it looks like this $1.00v + $1.80c = $98.80 So your two equations are: v+c=70 $1.00v + $1.80c = $98.80 So just rearrange the first equation so that it looks like this: v=70-c Now plug the rearranged equation into the second one find c. So you have $1.00(70-c) + $1.80c = $98.80 This gives us $70.00 - $1.00c +$1.80c = $98.80 Combine terms and rearrange gives us: .80c = $98.80-$70.00 So c = $18.80/.80 for an answer of 36 chicken wraps Now plug in the value of c (36) into the first equation to find the number of vegetarian wraps. So you have v + 36 wraps = 70 wraps So v= 70-36 Which gives us an answer of 34 vegetarian wraps.

OpenStudy (anonymous):

V + 1.8C = $98.8 V + C = 70 wraps Eliminate: -> .8 C = 28.8 => C = 28.8 / .8 = 36 chicken wraps Thus V = 70 - 36 = 34 vegetarian wraps.